# Data Mixing Can Induce Phase Transitions in Knowledge Acquisition

### v1

@@ 4.4 Data Mixing Induces Phase Transitions
To measure models’ performance on the first domain after training with algorithm $\mathcal{A}$ on the data mixture, we define ¯$\bar{\mathcal{L}}_{1}(\mathcal{A}):=\mathbb{E}_{\theta\sim\mathcal{P}_{1}}[\mathcal{L}(\mathcal{A}(\mathcal{D}_{\theta});\mathcal{D}_{\theta_{1}}^{(1)})]$ as the model’s expected loss on the first domain. If ¯$\bar{\mathcal{L}}_{1}(\mathcal{A})=F_{\mathcal{P}_{1}}(0)$, then the model learns nothing (random guessing). If ¯$\bar{\mathcal{L}}_{1}(\mathcal{A})=F_{\mathcal{P}_{1}}(\infty)$, the model perfectly learns the facts.
Theorem 4.3 shows that the learner sharply transitions between the two extremes as model size increases. This transition is characterized by two functions: $M^{-}_{0}(t):=\sup\{M\geq 0:-F^{\prime}_{\mathcal{P}_{2}}(M)>t\}$ and $M^{+}_{0}(t):=\inf\{M\geq 0:-F^{\prime}_{\mathcal{P}_{2}}(M)<t\}$. By rate-distortion theorem, $F_{\mathcal{P}_{2}}(M)$ is convex and hence $-F^{\prime}_{\mathcal{P}_{2}}(M)$ is non-increasing. Thus, $M^{-}_{0}(t)$ and $M^{+}_{0}(t)$ mark the last and first model sizes where $-F^{\prime}{\mathcal{P}_{2}}(M)$ exceeds or falls below $t$. If $F^{\prime}_{\mathcal{P}_{2}}(M)$ is strictly decreasing, then $M^{-}_{0}(t)=M^{+}_{0}(t)$.
@@
    if $M\leq M_{0}^{-}(\frac{r}{1-r}\cdot p)$, then ¯$\bar{\mathcal{L}}_{1}(\mathcal{A})=F_{\mathcal{P}_{1}}(0)$;
@@
    if $M\geq M_{0}^{+}(\frac{r}{1-r}\cdot p)+H_{\mathrm{tot}}$, then ¯$\bar{\mathcal{L}}_{1}(\mathcal{A})=F_{\mathcal{P}_{1}}(\infty)$.
@@
Consider the case where $F_{\mathcal{P}_{2}}(M)$ is a power-law function of $M$, i.e., $F_{\mathcal{P}_{2}}(M)=C+A\cdot M^{-\alpha}$. Here, $\alpha\in(0,1)$ and $A$ is a large constant. This is a reasonable assumption since LLM pre-training usually exhibits such power-law scaling behavior in model size  (Kaplan et al., 2020; Hoffmann et al., 2022). In this case, taking the derivative of $F_{\mathcal{P}_{2}}(M)$ gives $-F^{\prime}_{\mathcal{P}_{2}}(M)=A\cdot\alpha\cdot M^{-\alpha-1}$. Then, $M^{-}_{0}(t)=M^{+}_{0}(t)=(\frac{A\alpha}{t})^{1/(\alpha+1)}$. Plugging this into Theorem 4.3, we have the critical value for model size:
@@ 5 Power-Law Relationship of Threshold Frequency and Model Size
Figure 8: Validating the power-law relationship of threshold Frequency and model size. (a) & (b): Experiments on the mixture of SynBio-10k-power-law and FineWeb-Edu confirm that (1) the threshold frequency follows a power-law relationship with model size, and (2) the power-law exponent is approximately equal to the model scaling exponent plus one. (c) For the three open-source model families we examined, the threshold popularity for knowledge tested in PopQA also follows a power-law relationship with model size.
@@ 5.1 Experiments on Synthetic Biographies
As shown in Figure 8(a), $\log f_{\mathrm{thres}}$ and $\log M$ exhibit a linear relationship, yielding a slope of $1.152$. This value is larger than $1$, as expected from our theory. Further, we wonder if this slope is indeed close to $\alpha+1$. Following the approach of Hoffmann et al. (2022), we fit a model scaling function for FineWeb-Edu validation loss in Figure 8(b), obtaining $\alpha\approx 0.283$. This leads to a predicted exponent of $1.283$, which is close to the observed value of $1.152$.
@@ 5.2 Experiments on Knowledge Extracted from Wikipedia
We further evaluate models on PopQA (Mallen et al., 2023), which contains 14k QA pairs derived from Wikidata triplets, along with monthly page view for corresponding Wikipedia articles. Since knowledge tested in PopQA can be structured as triplets, we consider them as homogeneous and expect them to exhibit similar threshold frequencies for a given model size.
@@
Counting the frequency of specific knowledge in the pre-training data is challenging due to the scale (Kandpal et al., 2023). Following Mallen et al. (2023), we use Wikipedia page views as a proxy for popularity, which is assumed roughly proportional to the frequency of the knowledge in web data. To estimate the threshold popularity $P_{\mathrm{thres}}$, we identify the smallest popularity $P$ such that the model’s accuracy on knowledge with popularity above $P$ meets the target accuracy $\alpha_{\mathrm{target}}$ which is set to 60% in our experiments. See LABEL:sec:gpt-detail for details.
@@
We examine base models from Llama-2 (Touvron et al., 2023), Qwen-2.5 (Qwen et al., 2024), and Gemma-2 (Team et al., 2024), which are likely trained on similar data mixtures within each family. Figure 8(c) reveals that $\log P_{\mathrm{thres}}$ generally decreases linearly as $\log$ model size increases, though the slope varies across families due to differences in architecture and training data. We examine more model families in Section B.3.
@@ 6 Strategies to Enhance Knowledge Acquisition Under Low Mixing Ratios
Inspired by our theory, we propose two simple yet effective strategies to enhance knowledge acquisition under low mixing ratios—a common setting in practice, as a large $r$ may harm general capabilities expected to be acquired from multiple other data sources. The key idea is to raise the frequency of each fact, thereby increasing the “marginal value” of the knowledge-dense data.
@@
  • Strategy 2: Compact Knowledge Mixing (CKM): Rephrase the knowledge into a compact form and add the rephrased version to the original dataset while keeping the overall mixing ratio fixed.
We validate on both SynBio and a new real-world dataset, WikiBio, that these strategies significantly boost knowledge acquisition. For example, on WikiBio, subsampling and CKM improve the number of learned facts by 4 and 20 times, respectively. This is particularly surprising for subsampling, as it removes a significant proportion of the knowledge-dense data but ends up with higher accuracy.
@@ 6.1 Real-World Knowledge Data: WikiBio
The WikiBio Dataset.
To extend our study to a more real-world scenario, we curate WikiBio, a dataset containing Wikipedia biographies along with ten paraphrased versions of the first paragraph for 275k individuals, totaling 453M tokens. We ensure that the key information—name, occupation, and birth date—is mentioned within the first paragraph. This task is more challenging as Wikipedia biographies comprise diverse texts without uniform formats, requiring the model to generalize to prompts that rarely have exact matches in the training data. See Section C.2.2 for full details.
Evaluation.
We evaluate whether the model can recall a person’s birth date as a proxy for how well it memorizes the person’s information. Specifically, for a (name, occupation, birth date) triplet, we prompt the model with “The {occupation} {name} was born on” and consider the response correct if it includes correct the birth year and month in the generated text. The occupation is included not only to create out-of-distribution prompts but also to provide additional context.
@@ 6.2 Strategy 1: Random Subsampling
We study both pre-training from scratch and continual pre-training setups. To evaluate the model’s general capabilities, we use its validation loss on the web data (the Pile or FineWeb-Edu) and its zero-shot performance on five downstream tasks (see details in Section B.5). We compare the validation loss and average downstream performance to the model trained with $r=0$ in the pre-training-from-scratch setup or to the original Pythia model in the continual pre-training setup. Downstream performance drop of more than 2% is considered unacceptable.
@@
We train 410M models from scratch FineWeb-Edu mixed with SynBio-1.28M using $r\in\{0,0.1,0.2,0.3\}$ for a total of 32B tokens. As shown in Figures 9(a) and 11(a), increasing $r$ degrades FineWeb-Edu validation loss and downstream accuracy, with performance becoming unacceptable at $r=0.3$ (-2.09% accuracy, +0.05 loss), while SynBio accuracy remains near zero. In contrast, subsampling SynBio-1.28M to 25%, 50%, and 56.25% boosts SynBio accuracy to 23.53%, 37.46%, and 39.81%, respectively, while maintaining downstream performance within the acceptable range. Note that further increasing $\rho$ to 62.5% makes the frequency of each biography too low, resulting in SynBio accuracy dropping back to near zero. See more details in  LABEL:sec:mitigation_detail,  Figures 13(b) and 13(d).
@@
We continually pre-train the 410M and 1B Pythia models from their 100k-step checkpoints by mixing Pile with WikiBio and SynBio-2.56M, respectively. The 410M models are trained for 32B tokens and 1B models for 64B. Due to the distribution shift, the Pile validation loss may increase with training due to catastrophic forgetting (Ibrahim et al., 2024). To preserve models’ general capabilities, we apply early stopping when Pile validation loss increases by 0.05 (410M model) or 0.03 (1B model), each corresponding to $\sim 2\%$ drop in downstream performance. As shown in  Figures 9(b) and 11(b), without subsampling, $r=0.1$ or $0.15$ results in slow learning of WikiBio, while $r=0.2$ triggers early stopping after 20B tokens, resulting in poor WikiBio performance. By contrast, subsampling WikiBio to 25% or 50% significantly accelerates knowledge acquisition and keeps Pile validation loss acceptable. For example, for $r=0.1$, setting $\rho$ to 50% improves the number of learned facts by 4 times. Similar trends hold for 1B models: subsampling SynBio to 50% at $r=0.2$ outperforms both $r=0.2$ and early-stopped $r=0.4$ without subsampling by $\sim 30\%$. See more details in LABEL:sec:mitigation_detail, Tables 3, 13(e) and 13(c).
@@ 6.3 Strategy 2: Compact Knowledge Mixing (CKM)
The second strategy rephrases knowledge into compact forms (e.g., tuples) and adds them to the original dataset. Since the frequency of occurrence $f$ for each fact is inversely proportional to the average token count to represent each fact, adding compact representations can push $f$ above the threshold $f_{\mathrm{thres}}$ by reducing the average token count. For WikiBio, we compress the key information—name, birth date, and occupation—into a tuple format “Bio: N {name} B {birth date} O {occupation}”, and add these tuples until their token count reaches $\tau$ times the total token count of the original dataset. We call $\tau$ the CKM ratio.
@@
We apply CKM to WikiBio with the same continual pre-training setup as in  Section 6.2. Each time models encounter the tuple-form data point, the order of birth date and occupation is randomly flipped. We apply early stopping when Pile validation loss increases by 0.05.
@@
We explore CKM ratios $\tau\in\{0.1,0.3,0.6\}$ with $r=0.1$. As shown in Figures 9(b) and 11(c), CKM preserves the general capability and consistently boosts knowledge acquisition. Notably, performance on WikiBio improves by 4 times when $\tau$ is only 0.1. Increasing $\tau$ to 30% further boosts the number of learned facts by 20 times. See downstream performance in Table 4.
@@ 7 Discussions and Future Directions
This paper identifies phase transitions in knowledge acquisition under data mixing and provides information-theoretic explanations. While we mainly focus on factual knowledge acquisition, we also conduct preliminary experiments on a simple reasoning task. We emphasize that memorization is also important for reasoning—without basic knowledge, models cannot reason effectively, as also noted by Ruis et al. (2024); Xie et al. (2024). For example, solving math problems requires memorizing theorems, definitions, and techniques. We leave extensions to more complex reasoning tasks for future work.
@@
Following Allen-Zhu and Li (2024a), we use SynBio as a proxy for knowledge-dense data due to its clean and uniform format, which enables controlled experiments and easy measurement of knowledge acquisition. In contrast, real-world datasets are often messier and more heterogeneous—for example, Wikipedia includes both simple facts (e.g., biographies) and more complex content (e.g., scientific theories). These types of knowledge vary in learning difficulty and may exhibit different threshold frequencies. As a result, phase transitions may not be as apparent when mixing a heterogeneous dataset with web text. Nevertheless, our experiments on PopQA and WikiBio confirm that the theoretical insights still hold in real-world settings.
Appendix A Related Works

### v2

@@ 4.4 Data Mixing Induces Phase Transitions
To measure models’ performance on the first domain after training with algorithm $\mathcal{A}$ on the data mixture, we define $\bar{\mathcal{L}}_{1}(\mathcal{A}):=\mathbb{E}_{\theta\sim\mathcal{P}_{1}}[\mathcal{L}(\mathcal{A}(\mathcal{D}_{\theta});\mathcal{D}_{\theta_{1}}^{(1)})]$ as the model’s expected loss on the first domain. If $\bar{\mathcal{L}}_{1}(\mathcal{A})=F_{\mathcal{P}_{1}}(0)$, then the model learns nothing (random guessing). If $\bar{\mathcal{L}}_{1}(\mathcal{A})=F_{\mathcal{P}_{1}}(\infty)$, the model perfectly learns the facts.
Theorem 4.3 shows that the learner sharply transitions between the two extremes as model size increases. This transition is characterized by two functions: $M^{-}_{0}(t):=\sup\{M\geq 0:-F^{\prime}_{\mathcal{P}_{2}}(M)>t\}$ and $M^{+}_{0}(t):=\inf\{M\geq 0:-F^{\prime}_{\mathcal{P}_{2}}(M)<t\}$. By rate-distortion theorem, $F_{\mathcal{P}_{2}}(M)$ is convex and hence $F^{\prime}_{\mathcal{P}_{2}}(M)$ is non-increasing. Thus, $M^{-}_{0}(t)$ and $M^{+}_{0}(t)$ mark the last and first model sizes where $F^{\prime}{\mathcal{P}_{2}}(M)$ exceeds or falls below $t$. If $F^{\prime}_{\mathcal{P}_{2}}(M)$ is strictly decreasing, then $M^{-}_{0}(t)=M^{+}_{0}(t)$.
@@
    if $M\leq M_{0}^{-}(\frac{r}{1-r}\cdot p)$, then $\bar{\mathcal{L}}_{1}(\mathcal{A})=F_{\mathcal{P}_{1}}(0)$;
@@
    if $M\geq M_{0}^{+}(\frac{r}{1-r}\cdot p)+H_{\mathrm{tot}}$, then $\bar{\mathcal{L}}_{1}(\mathcal{A})=F_{\mathcal{P}_{1}}(\infty)$.
@@
Consider the case where $F_{\mathcal{P}_{2}}(M)$ is a power-law function of $M$, i.e., $F_{\mathcal{P}_{2}}(M)=C+A\cdot M^{-\alpha}$. Here, $\alpha\in(0,1)$ and $A$ is a large constant. This is a reasonable assumption since LLM pre-training usually exhibits such power-law scaling behavior in model size  (kaplan2020scaling; hoffmann2022an). In this case, taking the derivative of $F_{\mathcal{P}_{2}}(M)$ gives $F^{\prime}_{\mathcal{P}_{2}}(M)=A\cdot\alpha\cdot M^{-\alpha-1}$. Then, $M^{-}_{0}(t)=M^{+}_{0}(t)=(\frac{A\alpha}{t})^{1/(\alpha+1)}$. Plugging this into Theorem 4.3, we have the critical value for model size:
@@ 5 Power-Law Relationship of Threshold Frequency and Model Size
Figure 7: Validating the power-law relationship of threshold Frequency and model size. (a) & (b): Experiments on the mixture of SynBio-10k-power-law and FineWeb-Edu confirm that (1) the threshold frequency follows a power-law relationship with model size, and (2) the power-law exponent is approximately equal to the model scaling exponent plus one. (c): For the three open-source model families we examined, the threshold popularity for knowledge tested in PopQA also follows a power-law relationship with model size.
@@ 5.1 Experiments on Synthetic Biographies
As shown in Figure 7(a), $\log f_{\mathrm{thres}}$ and $\log M$ exhibit a linear relationship, yielding a slope of $1.152$. This value is larger than $1$, as expected from our theory. Further, we wonder if this slope is indeed close to $\alpha+1$. Following the approach of hoffmann2022an, we fit a model scaling function for FineWeb-Edu validation loss in Figure 7(b), obtaining $\alpha\approx 0.283$. This leads to a predicted exponent of $1.283$, which is close to the observed value of $1.152$.
@@ 5.2 Experiments on Knowledge Extracted from Wikipedia
We further evaluate models on PopQA (mallen-etal-2023-trust), which contains 14k QA pairs derived from Wikidata triplets, along with monthly page view for corresponding Wikipedia articles. Since knowledge tested in PopQA can be structured as triplets, we consider them as homogeneous and expect them to exhibit similar threshold frequencies for a given model size.
@@
Counting the frequency of specific knowledge in the pre-training data is challenging due to the scale (kandpal2023large). Following mallen-etal-2023-trust, we use Wikipedia page views as a proxy for popularity, which is assumed roughly proportional to the frequency of the knowledge in web data. To estimate the threshold popularity $P_{\mathrm{thres}}$, we identify the smallest popularity $P$ such that the model’s accuracy on knowledge with popularity above $P$ meets the target accuracy $\alpha_{\mathrm{target}}$ which is set to 60% in our experiments. See LABEL:sec:gpt-detail for details.
@@
We examine base models from Llama-2 (touvron2023llama), Qwen-2.5 (qwen2024qwen25technicalreport), and Gemma-2 (gemmateam2024gemma2improvingopen), which are likely trained on similar data mixtures within each family. Figure 7(c) reveals that $\log P_{\mathrm{thres}}$ generally decreases linearly as $\log$ model size increases, though the slope varies across families due to differences in architecture and training data. We examine more model families in Section D.4.
@@ 6 Strategies to Enhance Knowledge Acquisition Under Low Mixing Ratios
Inspired by our theory, we propose two simple yet effective strategies to enhance knowledge acquisition under low mixing ratios. This setting is common in practice, as a large $r$ may harm general capabilities expected to be acquired from other data sources. The key idea is to raise the frequency of each fact, thereby increasing the “marginal value” of the knowledge-dense data.
@@
  • Strategy 2: Compact Knowledge Mixing (CKM): Rephrase the knowledge into a compact form and add the rephrased version to the original dataset while keeping the overall mixing ratio fixed. See implementation details in LABEL:sec:mitigation_detail.
We validate these strategies on SynBio and WikiBio, a curated dataset of Wikipedia biographies. For example, on WikiBio, random subsampling and CKM improve the number of learned facts by 4x and 20x, respectively. The effectiveness of random subsampling is especially surprising, as it yields higher accuracy despite discarding a significant proportion of the knowledge-dense data.
@@ 6.1 Real-World Knowledge Data: WikiBio
To extend our study to a more real-world scenario, we curate WikiBio, a dataset containing Wikipedia biographies along with ten paraphrased versions for 275k individuals, totaling 453M tokens. This task is more challenging than SynBio as WikiBio features diverse texts without uniform formats, requiring the model to generalize to queries that rarely have exact matches in the training data. See Section E.2.2 for dataset construction details and LABEL:sec:mitigation_detail for evaluation details.
@@ 6.2 Strategy 1: Random Subsampling
We study both pre-training from scratch and continual pre-training setups. To evaluate the model’s general capabilities, we use its validation loss on the web data (the Pile or FineWeb-Edu) and its zero-shot performance on five downstream tasks (see details in Section D.7). We compare the validation loss and average downstream performance to the model trained with $r=0$ in the pre-training-from-scratch setup or to the original Pythia model in the continual pre-training setup. Downstream performance drop of more than 2% is considered unacceptable.
@@
We train 410M models from scratch FineWeb-Edu mixed with SynBio-1.28M using $r\in\{0,0.1,0.2,0.3\}$ for a total of 32B tokens. As shown in Figures 8(a) and 14(a), increasing $r$ degrades FineWeb-Edu validation loss and downstream accuracy, with performance becoming unacceptable at $r=0.3$ (-2.09% accuracy, +0.05 loss), while SynBio accuracy remains near zero. In contrast, subsampling SynBio-1.28M to 25%, 50%, and 56.25% boosts SynBio accuracy to 23.53%, 37.46%, and 39.81%, respectively, while maintaining downstream performance within the acceptable range. Note that further increasing $\rho$ to 62.5% makes the frequency of each biography too low, resulting in SynBio accuracy dropping back to near zero. See more details in  LABEL:sec:mitigation_detail,  Figures 15(b) and 15(d).
@@
We continually pre-train the 410M or 1B Pythia models from their 100k-step checkpoints on the mixture of the Pile and WikiBio or SynBio-2.56M. The 410M models are trained for 32B tokens and 1B models for 64B. When $r$ is large, the Pile validation loss may increase with training due to catastrophic forgetting (ibrahim2024simple). To preserve models’ general capabilities, we apply early stopping when Pile validation loss increases by 0.05 (410M model) or 0.03 (1B model), each corresponding to $\sim 2\%$ drop in downstream performance. As shown in  Figures 8(b) and 14(b), without subsampling, $r=0.1$ or $0.15$ results in slow learning of WikiBio, while $r=0.2$ triggers early stopping after 20B tokens, resulting in poor WikiBio performance. By contrast, subsampling WikiBio to 25% or 50% significantly accelerates knowledge acquisition and keeps Pile validation loss acceptable. For example, for $r=0.1$, setting $\rho$ to 50% improves the number of learned facts by 4 times. Similar trends hold for 1B models: subsampling SynBio to 50% at $r=0.2$ outperforms both $r=0.2$ and early-stopped $r=0.4$ without subsampling by $\sim 30\%$. See more details in LABEL:sec:mitigation_detail, Tables 4, 15(e) and 15(c).
@@ 6.3 Strategy 2: Compact Knowledge Mixing (CKM)
The second strategy rephrases knowledge into compact forms (e.g., tuples) and adds them to the original dataset. Given that the frequency of each fact $f$ is inversely proportional to its average representation token count, this augmentation reduces the average token count, thereby increasing $f$’s effective frequency and potentially pushing it above the threshold $f_{\mathrm{thres}}$. CKM is in the same spirit as the data synthesis technique in su2024nemotron, which rephrases high-quality data into condense forms such as QA pairs and knowledge lists.
For WikiBio, we compress the key information—name, birth date, and occupation—into the tuple format “Bio: N {name} B {birth date} O {occupation}”. We add these tuples until their token count equals a proportion $\tau$ (which we call the CKM ratio) of the original dataset’s token count.
@@
We apply CKM to WikiBio with the same continual pre-training setup as in  Section 6.2. We apply early stopping when Pile validation loss increases by 0.05.
@@
We fix $r=0.1$ and explore CKM ratios $\tau\in\{0.1,0.3,0.6\}$, which correspond to roughly 2x, 3x, and 4x increases in fact frequency, respectively. As shown in Figures 8(b) and 14(c), CKM preserves the general capability and consistently boosts knowledge acquisition. Notably, performance on WikiBio improves by 4x when the short-form augmentation makes up only 10% tokens of the original WikiBio dataset. Increasing $\tau$ to 30% further boosts the number of learned facts by 20x. See downstream performance in Table 5.
@@ 7 Discussions and Future Directions
Although our experiments mainly investigate factual knowledge, we also identify phase transitions in simple reasoning tasks. This suggests a commonality: memorization is foundational to reasoning, not just to fact-recall. Without basic knowledge, models cannot reason effectively, an observation shared by ruis2024procedural; xie2024memorization. For instance, solving math problems requires memorizing theorems, definitions, and techniques. We defer the exploration of more complex reasoning tasks to future work.
@@
Following allen2024physics, we use synthetic biographies as a proxy for knowledge-dense data for controlled and quantitative experiments. In contrast, real-world datasets are more heterogeneous—for example, Wikipedia includes both simple facts (e.g., biographies) and more complex content (e.g., scientific theories). These types of knowledge vary in learning difficulty and may exhibit different threshold frequencies. As a result, phase transitions may not be as apparent when mixing a heterogeneous dataset with web text. Nevertheless, our findings still apply at the level of individual knowledge pieces. That is, a specific fact or reasoning procedure may not be acquired at all if its frequency or the model size falls below a threshold, as evidenced by the theoretical result in (6) and empirical evidence in Section 5 and Section 3.3.
Acknowledgements and Disclosure of Funding
J.Z. acknowledges support by the National Key R&D Program of China 2024YFA1015800 and Shanghai Qi Zhi Institute Innovation Program.
Appendix A Limitations
The high computational costs to conduct all these experiments impede us from replicate all the experiments with different random seeds. These costs include the number of GPU hours. For example, a typical run of training a 410M model for 32B tokens requires 256 A100 GPU hours. Despite these difficulties, we managed to conduct experiments on models up to 6.9B and conduct ablation studies on hyperparameters in Section D.2.
Appendix B Broader Impact
This paper identifies two phase transitions in knowledge acquisition within data mixtures and provides theoretical understanding of these phenomena. Building on our theory, we propose two strategies to enhance the efficiency of knowledge acquisition. Our findings offer deeper insights into LLM behavior and can be applied to improve the factual accuracy of LLMs.
Appendix C Related Works